OpenStudy (lilsis76):

Can someone help me understand this and solve it please. Im really confused. The Reaction rate AB--> A + B is second order in AB and has a rate constant of 0.0118 M^-1 S^-1 at 25.0 Celsius. A reaction vessel initially contains 250.0 mL of 0.100 M AB which is allowed to react to form the gaseous product. The product is collected over water at 25.0 Celsius. How much time is required to produce 200.0 mL of the products at a barometric pressure of 755.1 mm Hg. (Vapor press. of H20 at this temp. is 23.8 mm Hg. )

OpenStudy (lilsis76):

@thomaster Do you think you can help me with this problem. We have covered so much in chemistry that the instructor gave us this problem and Im not sure what equation to use.

OpenStudy (aaronq):

i'm not entirely sure, but i think you have to first use the ideal gas law to find the moles of product, from the data in: " How much time is required to produce 200.0 mL of the products at a barometric pressure of 755.1 mm Hg. (Vapor press. of H20 at this temp. is 23.8 mm Hg." then you have to work with a form of the 2nd order-reaction equation to find the time required to reach the amount you found previously.

OpenStudy (lilsis76):



OpenStudy (aaronq):

use PV=nRT and find moles

OpenStudy (aaronq):

first.

OpenStudy (lilsis76):

okay. first Ill find the moles and I divided the.... hold on

OpenStudy (lilsis76):

250 mL x1000 = 2.5e 5L .100M / 2.5e5 L = 4.0e-7 moles right? @aaronq

OpenStudy (aaronq):

nope. use PV=nRT

OpenStudy (lilsis76):

errr.....okay, right now im trying to figure out how to fill it it. SO.....when im done can you tell me if it looks right before I proceed to solve it??? well, once i post it up

OpenStudy (lilsis76):

ok?! r is the moles right?? so far i have... @aaronq \[755.1mmHg(2.5x10^5L)=r(\frac{ 8.314 J }{ mol K })(298K)\] is this right? before I go on

OpenStudy (aaronq):

the volume part is wrong

OpenStudy (lilsis76):

okay so where the 2.5e5 L it should be 250.0 mL right?

OpenStudy (aaronq):

no i would use 200 mL or 0.2 L because you're finding the moles of product

OpenStudy (lilsis76):

200 mL, you divide it to get liters?

OpenStudy (lilsis76):

OH.........hold on haha

OpenStudy (lilsis76):

@aaronq \[755.1 mm Hg (.2 L) = r (\frac{ 8.314 J }{ mol K }) 298K\] once i do this i get r = .06095 J

OpenStudy (lilsis76):

YAY!!! @aaronq IM BACK!!! awww :( Ur off line now :'( Can someone else help me with this problem?? @thomaster

OpenStudy (lilsis76):

Site said it will be back on and the person that was helping me logged off. Can anyone of you please help me understand and work out this problem please. And thank you for taking the time to read this. @iambatman @thomaster @Compassionate @Euler271

OpenStudy (lilsis76):

@Jamierox4ev3r do you think u can help me understand and solve this please

HanAkoSolo (jamierox4ev3r):

yikes sorry :/

OpenStudy (lilsis76):

okay well So far aaronq was helping me before and told me to find the moles using the formula PV=nRT and I did. I rearranged it and got n= PV/RT giving me....\[\frac{ 755.1mmHG * .2 L }{ 8.314 * 298K } = n=.061 moles\]

OpenStudy (lilsis76):

@thomaster can you help me please? ?? @Jamierox4ev3r Okay, thank you anyways :/

OpenStudy (lilsis76):

@Euler271 do you think u can help me please? and Im sorry if it takes away for me to reply my computer keeps loading at times cuz there are alot of people on here.

OpenStudy (lilsis76):

@Data_LG2 Can u help me please Im superly stressing :(

OpenStudy (lilsis76):

YAY!!! UR BACK!!!, IM BACK!!! dumb computer website. It keeps losing connection and what not >:/

OpenStudy (aaronq):

sorry, the site crashed or something. you used the wrong R, and you didn't subtract the vapour pressure from the total. it should read: PV=nRT \((\dfrac{(755.1-23.8) mm Hg}{760 mmHg/atm})(0.2L)=n*(0.08602 \dfrac{L*atm}{K*mol})*298.15K\) ps. i converted to atm now you need to find the final concentration of AB. Can you figure that out? After use the 2nd order integrated equation: \(\dfrac{1}{[AB]_t}=\dfrac{1}{[AB]_0}+kt\) to find t, the time it takes.

OpenStudy (aaronq):

NOTE that when we found the moles, those were of A AND B, so take that into account.

OpenStudy (lilsis76):

Ya alot of people must be on here or something. okay then I found the right R= .0821 L*atm/ mol* K

OpenStudy (lilsis76):

okay. so moles...are A and B

OpenStudy (aaronq):

correct. you need to take that into consideration when finding moles of AB left, and subsequently, \([AB]_t\).

OpenStudy (lilsis76):

okay let me try again to find n for moles now that i have the right numbers

OpenStudy (lilsis76):

okay, I redid the calculations @aaronq and i got n=.0079 moles. Is this correct?

OpenStudy (aaronq):

i don't know, i'm not doing the calculations. i trust your arithmetic :)

OpenStudy (lilsis76):

oh haha okay. DUMB CONNNECTION!!! what is causing this?! it better not be my computer itself

OpenStudy (lilsis76):

any ways n=.0079 moles of AB.

OpenStudy (lilsis76):

To find the final concentration I use the 2nd order formula?

OpenStudy (aaronq):

nope, moles = 0.0079 of A + B and no, thats to find t once you know the final conc of AB

OpenStudy (lilsis76):

okay so .0079 are the moles of A + B so I have to divide it by 2 so each A, and B = .00395 moles

OpenStudy (lilsis76):

okay so..........................I have another question haha can u help me figure out the final concentraton of AB, I don't know how to do it after all. im just plugging in numbers and it doesn't look right.

OpenStudy (aaronq):

yeah 0.00395 moles is right, now use the volume of 250 mL, which i'm assuming is the volume of the vessel. Were assuming gas ideality in that the particles occupy a negligible amount of space.

OpenStudy (lilsis76):

okayim going to divide it and i get..... .00395/250=.0000158

OpenStudy (aaronq):

you need to watch your units 250 mL should be in L

OpenStudy (lilsis76):

Darn i! okay hold on let me try again. an I swithced computers I think my laptop is bad :/

OpenStudy (lilsis76):

okay i did the .00395/2.5x10^5 = 1.6x10^-8 M

OpenStudy (aaronq):

again you didn't convert correctly :P

OpenStudy (aaronq):

250 mL = 0.250 L

OpenStudy (lilsis76):

WHT? haha hold on let me check back Wait dont u multiply by 1000 to get to Liters?

OpenStudy (aaronq):

nope, divide

OpenStudy (lilsis76):

okay then, so when i go from L->mL i x1000 then

OpenStudy (lilsis76):

okay anyways.. .00395moles/.250L = .0158M

OpenStudy (aaronq):

yep, L to mL multiply by 1000 mL to L divide by 1000 okay, sweet. plug your values into the 2nd order integrated rate law

OpenStudy (lilsis76):

okay, another question haha [A]vt is inital right?

OpenStudy (lilsis76):